# Learning Hierarchical Sparse Transform Coding of 3DGS

### v2

@@ 4.1 Layer 1: Base Layer
Transform coding targets decorrelation and energy compaction; among orthonormal linear transforms, KLT is optimal for both. Under a tight parameter budget, we also find that an MLP-based transform lacks the capacity to surpass the KLT. Overall, the KLT offers an excellent cost–performance trade-off, achieving strong R-D performance with minimal compute and memory. We therefore adopt the KLT in Layer 1.
@@
Let $\mathbf{f}\in\mathbb{R}^{N_{f}}$ denote the anchor feature, with global mean $\mathbf{m}\in\mathbb{R}^{N_{f}}$ and KLT basis $\mathbf{V}\in\mathbb{R}^{N_{f}\times N_{f}}$. The analysis transform is $\boldsymbol{\theta}=g_{a}(\mathbf{f})=\mathbf{V}^{\top}(\mathbf{f}-\mathbf{m})$. The KLT coefficients exhibit an uneven energy distribution. For the anchor feature, we retain only the top-$M$ principal coefficients $\boldsymbol{\theta}_{p}=\boldsymbol{\theta}_{1:M}$, which are quantized and entropy-coded. The base layer reconstruction is $\hat{\mathbf{f}}_{base}=\mathbf{V}_{:,1:M}\hat{\boldsymbol{\theta}}_{p}+\mathbf{m}$.
4.2 Layer 2: Refinement Layer via Sparsity-Aware Residual Coding
Discarding the remaining coefficients causes information loss. The artifacts introduced by this truncation become more pronounced at high rates, leading to a quality bottleneck. However, coding all coefficients is rate-inefficient. To mitigate this information loss while keeping the rate overhead modest, we introduce a neural refinement layer. This layer transmits the KLT residual to compensate for truncation. Because most of the signal energy is captured by the leading KLT coefficients, the residual is dominated by low-magnitude components and can be well approximated as (approximately) sparse. For sparse or compressible signals, compressed sensing theory shows that the full signal can be reliably reconstructed from a small number of linear measurements by solving a sparsity-regularized optimization problem [donoho2006compressed]. This insight motivates us to use a principled, sparsity-aware design for the refinement transforms. Instead of relying on a generic multilayer perceptron (MLP), we adopt a parameter-efficient transform inspired by compressed sensing and deep unfolding techniques [zhang2020deep, zhang2018ista], enabling flexible compression of the residual with low parameter cost.
@@ 4.2.1 Analysis Transform: Learned Linear Measurements of Residuals
4.2.2 Synthesis Transform: Sparsity-Regularized Residual Reconstruction
@@
This inverse problem can be addressed using a variety of well-established optimization algorithms. A classical choice is the Iterative Shrinkage-Thresholding Algorithm (ISTA) [zibulevsky2010l1]. At the $k$-th iteration, ISTA updates the current estimate by performing a gradient descent step on the quadratic data fidelity term, followed by a soft-thresholding operation to handle the non-smooth $\ell_{1}$ regularization term. The update rule is given by:
@@
where $\eta$ denotes the step size, and $\mathcal{S}_{\tau}(\cdot)$ is the element-wise soft-thresholding function, defined as $\mathcal{S}_{\tau}(z)=\mathrm{sign}(z)\mathrm{max}(|z|-\tau,0)$. Directly applying ISTA, however, can be computationally expensive, as it typically requires many iterations and manual tuning of the hyperparameters $\tau$ and $\eta$. To address this, we adopt a deep unfolding approach [gregor2010learning, zhang2018ista, zhang2020deep] that maps a large number of ISTA iterations to a small number of learnable network layers, significantly reducing inference time while preserving interpretability and the sparsity prior. Each layer corresponds to one unfolded ISTA iteration, with layer-wise learnable step sizes and thresholds. In particular, $\tau$ and $\eta$ are learned separately for each channel in each layer. This design allows the synthesis transform to embed a learnable sparse structure, enabling efficient residual reconstruction from a compact representation with minimal overhead.
@@ 4.3 Implementation details
There are four types of objects that we need to compress:
  • Anchor positions: following HAC++ [chen2025hac++], we compress anchor coordinates using MPEG-GPCC [graziosi2020overview].
  • Features $\mathbf{f}\in\mathbb{R}^{50}$: we compress feature vectors with the full two-layer SHTC. The base layer encodes $\mathbf{f}$ into a low-dimensional representation $\boldsymbol{\theta}_{p}$ with $M=15$. The refinement layer then encodes the KLT residual $\mathbf{r}$ into a measurement vector $\mathbf{y}$ of dimension $N_{s}=15$.
  • Scalings $\mathbf{l}\in\mathbb{R}^{6}$: for the scaling vectors, we apply only the first SHTC layer (KLT) without dimensionality reduction. Because $\mathbf{l}$ is low-dimensional and the reconstruction quality is highly sensitive to it, truncating KLT coefficients in the transform domain severely degrades performance. Instead, for the KLT coefficients $\mathbf{l}_{t}\in\mathbb{R}^{6}$ we adopt a monotonically increasing, channel-wise quantization schedule:
    $q_{s}^{(i)}=q_{s}\exp(\alpha i),\quad i\in\{0,\ldots,5\},$ (8)
    where $\alpha$ is a learnable scalar and $q_{s}$ is the per-anchor base quantization step predicted by HAC. This quantization schedule acts as a remedial measure when dimensionality reduction cannot be applied: it assigns larger steps to lower-energy components, improving the rate–distortion trade-off while avoiding truncation loss.
  • Offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$: for offset vectors, we retain the original treatment in HAC. Due to offset masking, offset vectors have variable dimensionality across anchors, which makes a fixed transform impractical.
As described in Sec. 2 and illustrated in Fig. 1, HAC uses multi-scale hash grids to parameterize its entropy model. The original HAC queries multi-scale hash grids to predict the quantization step $q_{s}$ and Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$ for three types of anchor attributes: $\mathbf{f}$, $\mathbf{l}$, and $\{\mathbf{O}_{i}\}_{i=1}^{K}$. After integrating SHTC into HAC, we slightly modify the entropy coding process. It still queries multi-scale hash grids, but now predicts $q_{s}$ and $(\boldsymbol{\mu},\boldsymbol{\sigma})$ for the transformed vectors $\boldsymbol{\theta}_{p}$, $\mathbf{y}$, and $\mathbf{l}_{t}$, as well as for the offset vectors $\{\mathbf{O}_{i}\}_{i=1}^{K}$. The remaining quantization and coding procedures are the same as in the original HAC.
Inspired by color-space transforms in traditional codecs, we replace the original pixel-wise $\ell_{1}$ loss between the rendered image $\hat{I}$ and ground truth $I$ with a YCbCr-based loss, $\mathcal{L}_{\mathrm{YCbCr}}$, which computes reconstruction errors in the YCbCr color space. Because the human visual system is more sensitive to luminance than to chrominance variations, we assign a higher weight to the Y component and add a sharpness term on Y together with a smoothness term on Cb/Cr to preserve structure while suppressing color noise. This loss regularizes the compressed 3DGS representation, yielding fewer luminance errors and lower perceptual distortion at a given bitrate and thereby modestly improving R–D performance. The full definition of $\mathcal{L}_{\mathrm{YCbCr}}$ is provided in the supplementary material.
The overall training loss is defined as
[…]
5 Experimental results
5.1 Experimental setup
Dataset
We primarily evaluate the R-D performance on three commonly used large-scale real-scene datasets: DeepBlending [hedman2018deep], Mip-NeRF360 [barron2022mip], and Tanks&Temples [knapitsch2017tanks]. In particular, we assess all nine scenes from the Mip-NeRF360 dataset [barron2022mip]. Evaluations on additional datasets are provided in the supplementary material. These diverse datasets provide a comprehensive evaluation of the proposed SHTC framework.
Distortion metrics
We evaluate rendering distortion using three metrics: PSNR, SSIM [wang2004image], and LPIPS [zhang2018unreasonable].
Rate metrics
We adopt the BD-rate metric [bjontegaard2001calculation] to quantify memory savings by comparing the R–D curve of a compression method against that of a baseline. Specifically, BD-rate computes the average bitrate difference between the two R–D curves over a common distortion range. A negative BD-rate indicates that, on average, the method achieves a lower bitrate (measured in memory size) than the baseline at comparable distortion levels, with the absolute value (expressed as a percentage) quantifying the average bitrate reduction.
Training details
We follow the official training settings of HAC [chen2024hac] and HAC++ [chen2025hac++], and train each scene for 30k iterations. The Lagrange multiplier $\lambda$ is set to $\{0.002,0.004,0.008,0.015\}$ to obtain R–D curves that span a wide range of bitrates and distortions for a more comprehensive comparison. The gaps between successive $\lambda$ values are chosen to increase because the effect of $\lambda$ on the achieved bitrate is highly non-linear: especially in the low-bitrate regime, larger changes in $\lambda$ are required to produce clearly separated bitrates and thus more evenly distributed points on the R–D curve. We set $\lambda_{e}=0.03$ and $\lambda_{r}=\max(\lambda/4,0.001)$. All experiments are conducted on a server equipped with two RTX 4090 GPUs.
5.2 Comparison with Anchor-based Methods
5.2.1 Evaluating R-D performance
@@
Fig. 3 shows the R-D curves of several anchor-based 3DGS compression methods. Across all three datasets, the curve corresponding to the proposed SHTC, labeled “Ours SHTC (Full)” in the legend, lies toward the upper-left region, indicating higher PSNR at a given size or, equivalently, smaller size at a given distortion. For PCGS [chen2025pcgs], we train the model with $\lambda\in\{0.004,0.008,0.015,0.025\}$ to span a wide distortion range, so that its R-D points exhibit sufficient PSNR variation and a modest amount of overlap with the other methods, making the differences in the bitrate required to reach a given PSNR clearly visible. For HEMGS [liu2024hemgs], which is not open-sourced, we plot the R-D points reported in their paper and conduct an approximate comparison. Using only half as many training iterations (30k vs 60k), our method performs on par with or better than HEMGS on Mip-NeRF360 and DeepBlending and appears competitive on Tanks&Temples.
Tab. 1 reports the corresponding BD-rate of SHTC relative to HAC++, CAT-3DGS, and ContextGS, quantifying the bitrate (memory footprint) savings at equal PSNR. On average over the three datasets, our method achieves BD-rate values of about -21.0%, -24.3%, and -43.6% relative to HAC++, CAT-3DGS, and ContextGS, respectively.
Taken together, these results indicate that combining end-to-end optimized transforms with a simple entropy model is more effective than solely increasing entropy-model complexity in anchor-based 3DGS compression. This highlights the importance of incorporating transform coding and demonstrates the strong performance of the proposed SHTC framework.
Dataset HAC++ CAT-3DGS (40k) ContextGS Mip-NeRF360 -20.81% -24.54% -49.36% Tank&Temples -22.55% -22.78% -39.17% DeepBlending -19.58% -25.53% -42.30%
Table 1: The BD-rate of SHTC relative to three baselines: HAC++ [chen2025hac++], CAT-3DGS [zhan2025catdgs] and ContextGS [wang2024contextgs]. Negative values indicate bitrate savings (better) compared with each baseline.
5.2.2 Evaluating Computational Cost
@@
ContextGS HAC++ Ours ContextGS HAC++ Ours ContextGS HAC++ Ours ContextGS HAC++ Ours
Mip-NeRF360 106 128 146 32.73 8.48 5.90 33.32 13.86 8.86 -30.92% -56.60% -64.82%
Tank&Temples 130 158 177 29.37 5.63 3.91 29.05 8.81 5.80 -26.41% -44.87% -56.06%
DeepBlending 161 169 185 16.78 2.98 2.30 16.63 4.41 3.18 -41.09% -58.34% -64.56%
Table 2: Evaluating inference efficiency and compression effectiveness: complex entropy modeling alone vs. SHTC with a simple entropy model. From left to right, we report four metrics: rendering FPS, encoding time, decoding time, and BD-rate relative to the HAC baseline. All results are obtained from models trained with $\lambda=0.004$. The best result for each metric is highlighted in bold.
As shown in Tab. 2, our method not only achieves greater BD-rate gains relative to the HAC baseline than HAC++ and ContextGS, but also reduces both encoding and decoding latency. This efficiency gain stems from a paradigm shift: instead of relying on a highly complex entropy model to capture dependencies among high-dimensional anchor features, SHTC uses learned transforms to explicitly decorrelate anchor attributes before entropy coding. By shifting the decorrelation burden from the entropy model to the transform, SHTC produces coefficients for which a simple and fast entropy model is sufficient, yielding better overall compression efficiency and lower latency than relying on a highly complex entropy model alone.
In terms of training time, SHTC introduces a modest overhead. In our experiments, compared with HAC++, the training time per scene increases by at most about 8 minutes, while HAC++ itself already requires approximately 40–50 minutes to train. However, this additional cost is less critical in practice, because training is a one-time offline cost borne by 3DGS asset producers, whereas decoding is performed repeatedly on the user side: end users neither observe nor care about the training time, but they are directly sensitive to download and decoding latency. SHTC targets this more critical bottleneck by producing smaller bitstreams at comparable distortion and reducing decoding latency, making it particularly well suited for consumer-centric applications where fast access is essential. A detailed comparison of training times is provided in the supplementary material for reference.

### v3

@@ 3.3 SHTC: Sparsity-guided Hierarchical Transform Coding
Transform coding targets decorrelation and energy compaction; among orthonormal linear transforms, KLT is optimal for both. Under a tight parameter budget, we also observe that an MLP-based transform fails to outperform the KLT (see Appendix C). Overall, the KLT offers an excellent cost–performance trade-off, achieving strong R-D performance with minimal compute and memory. Therefore, we use the KLT as the foundation for constructing our transform. The KLT coefficients exhibit an uneven energy distribution. Discarding low-energy coefficients causes information loss. The artifacts introduced by this truncation become more pronounced at high rates, leading to a quality bottleneck. However, coding all coefficients is rate-inefficient. To balance these two extremes, we propose a hierarchical framework integrating KLT with a neural refinement layer.
@@
  • Layer 1: We use the KLT to construct a base layer for coarse reconstruction. Let $\mathbf{f}\in\mathbb{R}^{N_{f}}$ denote the anchor feature, with global mean $\mathbf{m}\in\mathbb{R}^{N_{f}}$ and KLT basis $\mathbf{V}\in\mathbb{R}^{N_{f}\times N_{f}}$. The base-layer analysis transform is given by $\boldsymbol{\theta}=g_{a}(\mathbf{f})=\mathbf{V}^{\top}(\mathbf{f}-\mathbf{m})$. For the anchor feature, we retain only the top-$M$ principal coefficients $\boldsymbol{\theta}_{p}=\boldsymbol{\theta}_{1:M}$, which are quantized and entropy-coded. The base-layer synthesis transform is given by $\hat{\mathbf{f}}_{base}=g_{s}(\hat{\boldsymbol{\theta}}_{p})=\mathbf{V}_{:,1:M}\hat{\boldsymbol{\theta}}_{p}+\mathbf{m}$.
  • Layer 2: To compensate for the truncation error of the base layer without incurring significant rate overhead, we introduce a neural refinement layer to transmit the KLT residual. Because most of the signal energy is captured by the leading KLT coefficients, the residual is is typically low-magnitude and approximately sparse/compressible. This observation raises a natural question: can we leverage this sparsity prior to introduce an inductive bias, enabling a very small refinement network to perform effective residual coding with strong R–D performance? We answer this question affirmatively by drawing on classical compressed sensing (CS), which suggests that a high-dimensional sparse (or compressible) signal can often be reconstructed from a small number of linear measurements via sparsity-regularized recovery (Donoho, 2006). Motivated by this insight, we design the refinement transforms in a sparsity-aware, CS-inspired manner. Specifically, the refinement-layer analysis transform produces a compact set of learned linear measurements of the residual, while the refinement-layer synthesis transform reconstructs the residual from the decoded measurements using an ISTA-style deep unfolding approach under a sparsity prior (Zhang and Ghanem, 2018; Zhang et al., 2020). A detailed description of these two transforms is provided next.
@@ 3.3.1 Analysis Transform: Learned Linear Measurements of Residuals
3.3.2 Synthesis Transform: Residual Reconstruction
@@
This inverse problem can be addressed using a variety of well-established optimization algorithms. A classical choice is the Iterative Shrinkage-Thresholding Algorithm (ISTA) (Zibulevsky and Elad, 2010). At the $k$-th iteration, ISTA updates the current estimate by performing a gradient descent step on the quadratic data fidelity term, followed by a soft-thresholding operation to handle the non-smooth $\ell_{1}$ regularization term. The update rule is given by:
@@
where $\eta$ denotes the step size, and $\mathcal{S}_{\tau}(\cdot)$ is the element-wise soft-thresholding function, defined as $\mathcal{S}_{\tau}(z)=\mathrm{sign}(z)\mathrm{max}(|z|-\tau,0)$. Directly applying ISTA, however, can be computationally expensive, as it typically requires many iterations and manual tuning of the hyperparameters $\tau$ and $\eta$. To address this, we adopt a deep unfolding approach (Gregor and LeCun, 2010; Zhang and Ghanem, 2018; Zhang et al., 2020) that maps a large number of ISTA iterations to a small number of learnable network layers, significantly reducing inference time while preserving interpretability and the sparsity prior. Each layer corresponds to one unfolded ISTA iteration, with layer-wise learnable step sizes and thresholds. In particular, $\tau$ and $\eta$ are learned separately for each channel in each layer. This design allows the synthesis transform to embed a learnable sparse structure, enabling efficient residual reconstruction from a compact representation with minimal overhead.
@@
Figure 5: Illustrative system pipeline integrating SHTC into HAC. Anchor coordinates are encoded by MPEG-GPCC, while SHTC latents and other attributes are quantized and entropy-coded using a hash-assisted context model that predicts probability parameters and quantization step sizes conditioned on $\mathbf{p}$.
@@ 3.4 Implementation details
We integrate the proposed SHTC into the HAC framework to build our 3DGS compressor. In Fig. 5, we illustrate how the SHTC analysis transform is integrated into the HAC pipeline before quantization and entropy coding. The synthesis transform mirrors the analysis transform and is omitted for brevity. Following HAC++ (Chen et al., 2025c), anchor coordinates $\mathbf{p}$ are compressed using MPEG-GPCC (Graziosi et al., 2020), yielding a separate coordinate bitstream. For the offset vectors $\{\mathbf{O}_{i}\}_{i=1}^{K}$, we keep the original HAC design. Due to offset masking, the number of valid offsets varies across anchors, making a fixed-dimensional transform impractical; therefore $\{\mathbf{O}_{i}\}_{i=1}^{K}$ are directly quantized and entropy coded as in HAC. For the scaling vector $\mathbf{l}$, we use only the base layer of SHTC (without dimensionality reduction) to produce $\mathbf{l}_{t}\in\mathbb{R}^{6}$, and then apply channel-wise step-size modulation during quantization following Eq.(8) with a dimension-adapted formulation, so that principal coefficients use smaller step sizes than less important ones; the resulting symbols are entropy-coded by HAC. For the feature vector $\mathbf{f}\in\mathbb{R}^{50}$, the full SHTC produces two latent representations, $\boldsymbol{\theta}_{p}$ and $\mathbf{y}$, with dimensions $M=15$ and $N_{s}=15$, respectively; both are quantized and entropy coded by HAC. As shown in Fig. 5, the probability parameters (e.g., mean and scale) and the quantization step $q_{s}$ for $\boldsymbol{\theta}_{p}$, $\mathbf{y}$, $\mathbf{l}_{t}$ and $\{\mathbf{O}_{i}\}_{i=1}^{K}$ are predicted from the multi-scale hash grids conditioned on the query position $\mathbf{p}$, and are used by the quantizer and the entropy coder.
We follow an HAC-style training objective and, in implementation, compute the pixel-wise reconstruction error in YCbCr space rather than RGB. Appendix E reports the full loss formulation and hyperparameter settings; the YCbCr distortion term is detailed in Appendix E.1.
@@
4 Experimental Results
4.1 Experimental Setup
@@
Dataset
We evaluate the R-D performance on five commonly used large-scale real-scene datasets: DeepBlending (Hedman et al., 2018), Mip-NeRF360 (Barron et al., 2022), Tanks&Temples (Knapitsch et al., 2017), Synthetic-NeRF (Mildenhall et al., 2021) and BungeeNeRF (Xiangli et al., 2022). In particular, we assess all nine scenes from the Mip-NeRF360 dataset (Barron et al., 2022). Evaluations on additional datasets are provided in the supplementary material. These diverse datasets provide a comprehensive evaluation of the proposed SHTC framework.
Distortion metrics
We evaluate rendering distortion using three metrics: PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018).
Rate metrics
Following prior work, we use the storage cost (measured as memory size in MB) as the rate metric. We adopt the BD-rate (Bjontegaard, 2001) to quantify rate savings by comparing the R-D curve of a method against a baseline. Specifically, BD-rate reports the average rate difference (in percentage) between two R-D curves over their overlapping distortion range. A negative BD-rate indicates that the method requires less storage than the baseline at comparable distortion levels, and the magnitude reflects the average percentage of rate reduction.
4.2 Comparison with Anchor-based Methods
4.2.1 Evaluating R-D performance
Figure 6 shows the R-D curves of several anchor-based 3DGS compression methods. Across all three datasets, the curve corresponding to the proposed SHTC, labeled “Ours SHTC (Full)” in the legend, lies toward the upper-left region, indicating higher PSNR at a given size or, equivalently, smaller size at a given distortion. For HEMGS (Liu et al., 2024a), which is not open-sourced, we plot the R-D points reported in their paper and conduct an approximate comparison. Using only half the training iterations of HEMGS (30k vs. 60k), our method performs on par with or better than HEMGS. The R-D curves tested on Synthetic-NeRF and BungeeNeRF are provided in Fig. 7 in Appendix. Table 1 reports the corresponding BD-rate of our method (HAC + Ours SHTC) relative to HAC++, CAT-3DGS, ContextGS and HAC, quantifying the average memory footprint savings at equal PSNR. These results demonstrate the superiority of our proposed TTC paradigm over anchor-based compression, which relies solely on increasingly complex entropy models.
Note that SHTC is not tailored to HAC and does not depend on it. When integrated into ContextGS, SHTC also yields consistent improvements over the vanilla ContextGS baseline. The R-D curves are shown in Fig. 8 in Appendix, and ContextGS+SHTC achieves a BD-rate of $19.45\%$ relative to vanilla ContextGS.
Table 1: The BD-rate of our method relative to four baselines: HAC++, CAT-3DGS, ContextGS and vanilla HAC. Negative values indicate bitrate savings (better) compared with each baseline.
Dataset HAC++ CAT-3DGS (40k) ContextGS HAC Mip-NeRF360 -20.81% -24.54% -49.36% -64.82% Tank&Temples -22.55% -22.78% -39.17% -56.06% DeepBlending -19.58% -25.53% -42.30% -64.56% BungeeNeRF -10.04% -39.90% - -52.05% Synthetic-NeRF -13.45% - - -26.33%
Table 2: Efficiency and rate–distortion comparison on Mip-NeRF360 ($\lambda=0.004$), evaluated on the same RTX 4090 server. HAC is reported as a reference anchor and is not included in the comparison. We compare Ours against ContextGS, CAT-3DGS, and HAC++ in terms of training time (in seconds), rendering FPS, encoding/decoding latency (in seconds), and BD-rate measured relative to the HAC baseline. Best results among the compared methods are highlighted in bold.
Methods Training Rendering Encoding Decoding BD-rate Time FPS Time Time HAC 1949 131 4.20 10.05 0 ContextGS 3927 106 32.73 33.32 -30.92% CAT-3DGS 5725 - 79.82 87.95 -53.48% HAC++ 2735 128 8.48 13.86 -56.60% Ours 3109 146 5.85 8.84 -64.82%
4.2.2 Evaluating Transform Overhead
Computational overhead.
We use vanilla HAC as the reference baseline and compare our method with state-of-the-art anchor-based codecs, including HAC++, ContextGS, and CAT-3DGS. We report BD-rate relative to HAC to quantify R-D gains, and we additionally measure encoding/decoding latency and training time to characterize the cost of these improvements. Results are summarized in Table 2 and visualized in Fig. 2. Overall, our method achieves improved R–D performance while maintaining low coding latency. We attribute this advantage to the introduced transforms, which produce features that better match a lightweight, parallel-friendly entropy model, thereby avoiding expensive and complex entropy modeling. In terms of training cost, SHTC incurs substantially less overhead than CAT-3DGS and ContextGS, while being moderately slower than HAC++. Since many 3DGS deployments follow a “compress once, decompress many times” workflow, this training overhead is a one-time offline cost for asset producers and typically does not affect client-side user experience. In contrast, the memory footprint and decoding latency directly impact client-side user experience and are therefore the primary practical constraints. Consequently, the proposed SHTC method offers a favorable trade-off and lies on (or close to) the empirical Pareto frontier among the compared anchor-based methods as shown in Fig. 2.
Parameter overhead.